### v2

@@ 6.2 Binary-action supermodular games
Let $(T,\Theta,\pi,u)$ be a supermodular game with $A_{i}=\{0,1\}$ and $\Theta=\Theta^{*}$. For $S\subseteq I$, we denote by $\mathbf{1}_{S}$ the action profile in which all players in $S$ play action 1, while the others play action 0. By convention, we write $\mathbf{1}=\mathbf{1}_{I}$. We assume that this game has a generalized potential function $F:\mathcal{A}\times\Theta\to\mathbb{R}$ satisfying the following conditions:
@@
By Theorem 1, $\sigma^{\mathbf{1}}$ is robust for arbitrary prior $\pi\in\Delta(T\times\Theta)$. Moreover, the converse is also true: if $\sigma^{\mathbf{1}}$ is a robust BIBCE of a binary-action supermodular game for arbitrary prior $\pi\in\Delta(T\times\Theta)$, then it must (generically) be a GP-maximizing BIBCE with a genelized potential function satisfying G1 and G2, as formalized in the next proposition.
@@
$v(\mathbf{1},\theta)>v(a,\theta)\text{ for all }a\neq\mathbf{1}\text{ and }\theta\in\Theta.$ (22)
@@
Oyama and Takahashi (2020) focus on complete information games and show that, for a generic binary-action supermodular game, if there does not exist a monotone potential function satisfying (22), then, for every $\varepsilon>0$, there exists an $\varepsilon$-elaboration in the sense of Kajii and Morris (1997) such that “all 00” survives the iterated elimination of strictly dominated strategies. Since an $\varepsilon$-elaborations in Kajii and Morris (1997) is equivalent to our $\varepsilon$-elaboration (see Subsection 7.1), the action profile $\mathbf{1}$ is not robust in our sense.
This result leads us to part (2) of Proposition 2. Suppose there is no generalized potential function satisfying G1 and G2. Then, there cannot be a monotone potential function satisfying (22). This, in turn, means that there exists $\theta^{\prime}\in\Theta$ such that the complete information game in which $\theta^{\prime}$ is common knowledge does not have a monotone potential function satisfying (22) with $\theta=\theta^{\prime}$. Now, consider a degenerate prior $\pi^{\prime}\in\Delta(T\times\Theta)$ with $\pi^{\prime}(t^{\prime},\theta^{\prime})=1$ for some specific type $t^{\prime}\in T$. Since $(T,\Theta,\pi^{\prime},u)$ is a complete information game, there exists an $\varepsilon$-elaboration in which “all 00” survives the iterated elimination of strictly dominated strategies. This establishes that $\sigma^{\mathbf{1}}$ is not robust in $(T,\Theta,\pi^{\prime},u)$.
@@ B.3 Proof of Lemma 7
$B^{\mathcal{A}}(\sigma)=\{\sigma^{\prime}\in\Sigma^{\mathcal{A}}\mid\text{$\sigma_{i}^{\prime}$ is a best response strategy to $\sigma_{-i}$ for each $i\in I$}\}$
@@ C.1 Proof of Lemma 8
$B^{k,\mathcal{A}}(\sigma)=\{\sigma^{\prime}\in\Sigma^{k,\mathcal{A}}\mid\text{$\sigma_{i}^{\prime}$ is a best response strategy to $\sigma_{-i}$ for each $i\in I$ of $t_{i}\in T^{k}$}\}$
@@
[…]

### v4

@@ 6.2 Binary-action supermodular games
Let $(T,\Theta,\pi,u)$ be a supermodular game with $A_{i}=\{0,1\}$ and $\Theta=\Theta^{*}$. For $S\subseteq I$, we denote by $\mathbf{1}_{S}$ the action profile in which all players in $S$ choose action 1, while the others choose action 0. By convention, we write $\mathbf{1}=\mathbf{1}_{I}$. We assume that this game has a generalized potential function $F:\mathcal{A}\times\Theta\to\mathbb{R}$ satisfying the following conditions:
@@
By Theorem 1, $\sigma^{\mathbf{1}}$ is robust for arbitrary prior $\pi\in\Delta(T\times\Theta)$. Moreover, the converse is also true: if $\sigma^{\mathbf{1}}$ is a robust BIBCE of a binary-action supermodular game for arbitrary prior $\pi\in\Delta(T\times\Theta)$, then it must (generically) be a GP-maximizing BIBCE with a generalized potential function satisfying G1 and G2, as formalized in the next proposition.
@@
[…]
@@
Oyama and Takahashi (2020) focus on complete information games and show that, for a generic binary-action supermodular game, if there does not exist a monotone potential function satisfying (22), then, for every $\varepsilon>0$, there exists an $\varepsilon$-elaboration in the sense of Kajii and Morris (1997) such that “all 0” survives the iterated elimination of strictly dominated strategies. Since an $\varepsilon$-elaborations in Kajii and Morris (1997) is equivalent to our $\varepsilon$-elaboration (see Subsection 7.1), the action profile $\mathbf{1}$ is not robust in our sense.
This result leads us to part (2) of Proposition 2. Suppose there is no generalized potential function satisfying G1 and G2. Then, there cannot be a monotone potential function satisfying (22). This, in turn, means that there exists $\theta^{\prime}\in\Theta$ such that the complete information game in which $\theta^{\prime}$ is common knowledge does not have a monotone potential function satisfying (22) with $\theta=\theta^{\prime}$. Now, consider a degenerate prior $\pi^{\prime}\in\Delta(T\times\Theta)$ with $\pi^{\prime}(t^{\prime},\theta^{\prime})=1$ for some specific type $t^{\prime}\in T$. Since $(T,\Theta,\pi^{\prime},u)$ is a complete information game, there exists an $\varepsilon$-elaboration in which “all 0” survives the iterated elimination of strictly dominated strategies. This establishes that $\sigma^{\mathbf{1}}$ is not robust in $(T,\Theta,\pi^{\prime},u)$.
@@ B.3 Proof of Lemma 7
[…]
@@ C.1 Proof of Lemma 8
[…]
@@
$\text{ where $\gamma^{1}\in\Psi^{1}(\gamma)$ and $\ \gamma^{2}\in\Psi^{2}(\gamma)$}\}.$